OpenStudy (anonymous):

Passing through (-3,3) and parallel to the line whose equation is 3x-2y=5? need point slope intercept form

OpenStudy (anonymous):

Point slope intercept form? Which one? Point-slope or slope-intercept form?

OpenStudy (anonymous):

slope intercept form

OpenStudy (anonymous):

i got the poin slope but i'm having trouble with the slope intercept form for some reason.

OpenStudy (anonymous):

Just plug in the x and y into this: 3x - 2y = 3(-3) - 2(3) = -15 3x - 2y = -15 Now, just solve for y. 3x - 2y = -15 -2y = -3x - 15 y = 1.5x + 7.5

OpenStudy (anonymous):

okay will do thanks

OpenStudy (anonymous):

No problem :)